1.) y=4x 3x+y=-21
2.) y=2x-2 2x+3y=10

Answer by mananth(16949)

y=4x 3x+y=-21
y-4x=0 (1)
y+3x=-21 (2)by rewriting the equation
substract (2) from (1)
Change the signs of the eq. No 2 and add the common terms
-7x=-21
-x=-21/7
-x=-7 or x=7


2.) y=2x-2 2x+3y=10
y-2x=-2 (1)
3y+2x=10 (2)
Add the two equations
4y=8
so y =8/4 =2 y=2